# Double robust estimation of functional outcomes with data missing at random

### v2

@@ Abstract
We present and study semi-parametric estimators for the mean of functional outcomes in situations where some of these outcomes are missing and covariate information is available on all units. Assuming that the missingness mechanism depends only on the covariates (missing at random assumption), we present two estimators for the functional mean parameter, using working models for the functional outcome given the covariates, and the probability of missingness given the covariates. We contribute by establishing that both these estimators have Gaussian processes as limiting distributions and explicitly give their covariance functions. One of the estimators is double robust in the sense that the limiting distribution holds whenever at least one of the nuisance models is correctly specified. These results allow us to present simultaneous confidence bands for the mean function with asymptotically guaranteed coverage. A Monte Carlo study shows the finite sample properties of the proposed functional estimators and their associated simultaneous inference. The use of the method is illustrated in an application where the mean of counterfactual outcomes is targeted.
Causal inference,
functional data analysis,
lifetime income trajectory,
simultaneous confidence bands,
keywords:
[…]
, , , and
@@ 1 Introduction
We propose and study semi-parametric estimators of the mean of functional outcomes when these functions are not observed for all individuals in the study. Functional outcomes arise in many different situations, e.g. when studying human movement (Dannenmaier et al., 2020) or lifetime cumulative income trajectories (Ecker et al., 2022), (see also Ullah and Finch, 2013, for further examples). Missingness of the outcomes can occur due to e.g., non-response (in a survey context) or drop-out (for follow-up studies). Missingness can also be a consequence of considering the potential outcome framework for causal inference allowing us to answer counterfactual questions; see Holland (1986), Ding and Li (2018), and Section 4 for an application. We focus here on the case where some baseline covariates are observed for all units, and the missingness mechanism is governed only by these covariates. This is referred to as the missing at random assumption (see e.g., Tsiatis, 2006; Little and Rubin, 2019). Because the entire functional outcome is missing for some units, our context is different from the literature on fragmented functional data, where it is assumed that functional objects may only be observed on some subset(s) of their domain. This literature does typically not consider access to covariates and instead aims to reconstruct the missing sections using only information from the observed functional outcomes (Delaigle and Hall, 2016; Kraus, 2019; Kneip and Liebl, 2020).
There is a large body of literature on semi-parametric estimators of the mean when a scalar outcome is missing at random (e.g., Tsiatis, 2006; Little and Rubin, 2019; Molenberghs et al., 2015). Popular estimators include outcome regression (also called regression imputation) estimators (e.g., Tan, 2007; Genbäck and de Luna, 2019) and augmented inverse probability weighted estimators (also called double robust estimators; e.g., Robins et al., 1994; Bang and Robins, 2005; Kang and Schafer, 2007). The outcome regression (OR) method involves modelling the relationship between the outcome and the observed covariates and imputing the missing outcomes with predictions from this model. The double robust (DR) estimator can be interpreted as a correction of the OR estimator, using weights that are based on modelling the missingness mechanism as a function of the covariates. The DR estimator is locally efficient (it attains the semi-parametric efficiency bound if both models are correct) and double robust (consistent and asymptotically normal if only one of the two models is correct; (e.g., Seaman and Vansteelandt, 2018)).
@@
These asymptotic properties allow us to use recent results in Liebl and Reimherr (2023) to propose simultaneous confidence bands for the functional mean, providing a given coverage level over the entire domain. At the same time, these bands also allow for local inferences and the interpretation of local significance levels over user-determined partitions of the domain. Up to our knowledge, there is only one earlier attempt to study semi-parametric estimators in the present context. Belloni et al. (2017), in the causal inference context, study double robust estimators and obtain their limiting distribution, but without an expression for the covariance function. Instead, they propose a resampling scheme to obtain valid inference. In another related contribution, Ecker et al. (2024) study an OR estimator for a functional causal effect and provides its finite sample distribution, albeit under stricter distributional assumptions about the underlying data.
@@ 2.1 Model and semi-parametric estimation
Suppose we have a random sample, $(\mathcal{Y}_{i},\textbf{x}_{i},Z_{i})_{i=1}^{n}$, from the joint distribution of $(\mathcal{Y},\textbf{x},Z)$, where the outcome variable, $\mathcal{Y}$, is a $L^{2}([0,1],\mathbb{R})$ valued random element, and the set of covariates, x, is a $p\times 1$ random vector with mean vector $\boldsymbol{\mu}_{x}$ and covariance matrix $\boldsymbol{\Sigma}_{x}$. Further, $Z$ is a binary variable indicating whether the outcome $\mathcal{Y}$ is observed, in which case $Z=1$, or missing ($Z=0$). We assume throughout that $\Pr[Z_{i}=1|\textbf{x}_{i},\mathcal{Y}_{i}]=\Pr[Z_{i}=1|\textbf{x}_{i}]$, that is, the missingness mechanism is ignorable (also referred to as the outcome being missing at random, (Tsiatis, 2006, Chap 6)). The parameter of interest is the functional mean $\mu_{y}=\mathbb{E}\left(\mathcal{Y}\right)$.
@@
$\mathcal{Y}_{i}=\textbf{x}_{i}^{\top}\boldsymbol{\beta}+\epsilon_{i},$ (1)
where $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{p})$ and $\beta_{j}$s are non-stochastic functions in $L^{2}([0,1],\mathbb{R})$, and $\epsilon_{i}$ is a $L^{2}([0,1],\mathbb{R})$ valued random element with zero mean and with $\sigma_{\epsilon}$ as its covariance function. Furthermore, a working model for $\Pr[Z_{i}=1|\textbf{x}_{i}]$ is specified as the logistic regression with probability of success:
$\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}),$ (2)
@@
$\hat{\mu}_{OR}=n^{-1}\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}$, (3)
where $\hat{\boldsymbol{\beta}}$ is the functional estimation $\left(\mathbf{X}^{\top}\mathbf{X}\right)^{-1}\mathbf{X}^{\top}\boldsymbol{\mathcal{Y}},$ with $\boldsymbol{\mathcal{Y}}=\left(\mathcal{Y}_{1},\dots,\mathcal{Y}_{n}\right)^{\top}$, and $\mathbf{X}=\left(Z_{1}\mathbf{x}_{1},\dots,Z_{n}\mathbf{x}_{n}\right)^{\top}$. Based on the two working models (1) and (2), the double robust estimator of ${\mu}_{y}$, is defined as
@@
where $\hat{\boldsymbol{\gamma}}$ is the maximum likelihood estimator of $\boldsymbol{\gamma}$.
@@ 2.2 Main results
Our strategy to obtain the asymptotic distribution of the estimators (3) and (4) consists of two steps that allow us to apply Prohorov’s Theorem (see Billingsley, 1979, Sect. 5, p.57). First, we consider arbitrary finite-dimensional distributions (i.e. based on a finite discretization on [0,1]) and deduce their joint asymptotic distributions (Lemma A.1 and A.2 in the Appendix). These results are the consequence of the theory on M-estimators (Newey and McFadden, 1994). We provide in these lemmas an explicit characterization of the asymptotic covariance function of multivariate OR and DR estimators, which we cannot find in the earlier literature. In a second step, given the finite-dimensional results, the functional asymptotic results are established by demonstrating tightness. For this purpose, the re-scaled OR estimator, $\sqrt{n}\left(\hat{\mu}_{OR}-\mu_{y}\right)$, can be rewritten as:
$\sqrt{n}\left(\bar{\textbf{x}}_{n}-\boldsymbol{\mu}_{x}\right)^{\top}\boldsymbol{\beta}+\left(\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\right)\left(\sum_{i=1}^{n}Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right)^{-1}\boldsymbol{\zeta}.$ (5)
where $\boldsymbol{\zeta}=\left(n^{-1/2}\sum_{i=1}^{n}{Z_{i}x_{i,1}\epsilon_{i}},\dots,n^{-1/2}\sum_{i=1}^{n}{Z_{i}x_{i,p}\epsilon_{i}}\right)^{\top}$. Similarly, the re-scaled DR estimator, $\sqrt{n}\left(\hat{\mu}_{DR}-\mu_{y}\right)$, can be represented as
$\sqrt{n}\left(\hat{\mu}_{OR}-\mu_{y}\right)-\left(\sum_{i=1}^{n}\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\boldsymbol{\gamma}})}\textbf{x}_{i}^{\top}\right)\left(\sum_{i=1}^{n}Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right)^{-1}\boldsymbol{\zeta}+n^{-1/2}\sum_{i=1}^{n}\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\boldsymbol{\gamma}})}\epsilon_{i}.$ (6)
@@
It is woth noting that the functional limiting distribution of the OR estimator (3) could be obtained by applying functional asymptotic theorems developed in recent years, e.g. Telschow and Schwartzman (2022). However, the functional DR estimator (4) presents more of a challenge, therefore the strategy outlined above is required. To maintain consistency throughout the paper, this strategy is applied to address both estimators (3) and (4). This allows us to obtain the following results on the asymptotic distributions for these functional estimators.
Theorem 2.3.
Assume that the working model (1) is correctly specified, i.e., $E(\mathcal{Y}_{i}\mid\textbf{x}_{i})=\textbf{x}_{i}^{\top}\boldsymbol{\beta}$. Then the functional OR estimator (3) is asymptotically distributed as a Gaussian process:
$\sqrt{n}(\hat{\mu}_{OR}-\mu_{y})\overset{d}{\rightarrow}\mathcal{GP}(0,\boldsymbol{\beta}(s)^{\top}\boldsymbol{\Sigma}_{\textbf{x}}\boldsymbol{\beta}(t)+\boldsymbol{\mu}_{\textbf{x}}^{\top}\boldsymbol{\Pi}^{-1}\boldsymbol{\mu}_{\textbf{x}}\cdot\sigma_{\epsilon}(s,t)).$
@@
By Prohorov’s Theorem, this theorem is a consequence of the results in Lemma A.2, 2.1, and 2.2; see Appendix A and B for details. ∎
Theorem 2.4.
Assume that at least one of the working models (1) and (2) is correctly specified, i.e., $E(\mathcal{Y}_{i}\mid\textbf{x}_{i})=\textbf{x}_{i}^{\top}\boldsymbol{\beta}$ or $\Pr[Z_{i}=1|\textbf{x}_{i}]=\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})$. Then, the functional DR estimator $\sqrt{n}(\hat{\mu}_{DR}-\mu_{y})$ is asymptotically distributed as a Gaussian process with zero mean function. Further, if both working models are correctly specified, the covariance function simplifies and we have:
$\sqrt{n}(\hat{\mu}_{DR}-\mu_{y})\overset{d}{\rightarrow}\mathcal{GP}(0,\boldsymbol{\beta}(s)^{\top}\boldsymbol{\Sigma}_{\textbf{x}}\boldsymbol{\beta}(t)+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})]\sigma_{\epsilon}(s,t)).$
@@
By Prohorov’s Theorem, this theorem is a consequence of the results in Lemma A.2, 2.1, and 2.2; see Appendix A and B for details. ∎
Note that Theorem 2.4 yields a double robustness property for the functional estimator, whereby it is asymptotically distributed as a Gaussian process with zero mean function (concentrates at root-$n$ rate) when only one of the nuisance models is correctly specified. Given that we have functional estimators with Gaussian processes as their limiting distribution, we can obtain simultaneous confidence bands as described below.
@@ 2.3 Feasible estimation and simultaneous confidence bands
The functional estimators (3) and (4) are not directly feasible since the functional outcomes are infinite dimensional objects. However, considering functional estimators and their derived asymptotic distribution allows us to perform simultaneous inference on feasible counterparts. An example of feasible estimators that we use here are the finite-dimensional estimators based on a discretization of the domain in (A.2) and (A.3) in Appendix A.
Simultaneous confidence bands for the functional parameter of interest, $\mu_{y}$, can be constructed by applying the results obtained by Liebl and Reimherr (2023). This work introduces ”fast and fair” simultaneous confidence bands based on the construction of an adaptive, non-constant critical value function, which ensures that false positive rates are balanced across different partitions of the domain. The method enables both global and local interpretations of the bands and is less computationally expensive compared to many simulation-based and resampling-based approaches (e.g., Pini and Vantini, 2017; Abramowicz et al., 2018).
To apply the results in Liebl and Reimherr (2023), we need to further restrict $\mathcal{Y}_{i}$ to be once continuously differentiable almost surely in the sequel. Under this assumption and the conditions of Theorem 2.3, a valid 95% simultaneous confidence band (SCB) for $\mu_{y}$ based on the OR estimator is given by:
@@
where $C_{OR}$ is the covariance function of the OR estimator given in Theorem 2.3, and $\mathit{u}^{\star}_{2.5}(t)$ is a critical value function chosen such that
@@
Simultaneous confidence bands for $\mu_{y}$ based on the DR estimator can be constructed analogously by using $\hat{\mu}_{DR}$ and $C_{DR}$ given in Theorem 2.4. See Liebl and Reimherr (2020, 2023) for further details on the construction and properties of these simultaneous confidence bands.

### v3

# Doubly robust estimation with functional outcomes missing at random

@@ Abstract
We present and study semi-parametric estimators for the mean of functional outcomes in situations where some of these outcomes are missing, and covariate information is available on all units. Assuming that the missingness mechanism depends only on the covariates (missing at random assumption), we present two estimators for the functional mean parameter, using working models for the functional outcome given the covariates, and the probability of missingness given the covariates. We contribute by establishing that both these estimators have Gaussian processes as limiting distributions and explicitly give their covariance functions. One of the estimators is doubly robust in the sense that the limiting distribution holds whenever at least one of the nuisance models is correctly specified. These results allow us to present simultaneous confidence bands for the mean function with asymptotically guaranteed coverage. A Monte Carlo study shows the finite sample properties of the proposed functional estimators and their associated simultaneous inference. The use of the method is illustrated in an application where the mean of counterfactual outcomes is targeted.
Keywords: Causal inference; functional data analysis; lifetime income trajectory; simultaneous confidence bands
@@ 1 Introduction
We propose and study semi-parametric estimators of the mean of functional outcomes when these functions are not observed for all individuals in the study. Functional outcomes arise in many different situations, e.g. when studying human movement (dannenmaier2020application, ) or lifetime cumulative income trajectories (ecker2022regional, ), (see also ullah2013applications, , for further examples). Missingness of the outcomes can occur due to e.g., non-response (in a survey context) or drop-out (for follow-up studies). Missingness can also be a consequence of considering the potential outcome framework for causal inference allowing us to answer counterfactual questions; see holland86 , missingcausal , and Section 4 for an application. We focus here on the case where some baseline covariates are observed for all units, and the missingness mechanism is governed only by these covariates. This is referred to as the missing at random assumption (see e.g., tsiatis2006semiparametric, ; little2019statistical, ). Because the entire functional outcome is missing for some units, our context is different from the literature on fragmented functional data, where it is assumed that functional objects may only be observed on some subset(s) of their domain. This literature does typically not consider access to covariates and instead aims to reconstruct the missing sections using only information from the observed functional outcomes delaigle2016approximating ; kraus2019inferential ; kneip2020optimal .
There is a large body of literature on semi-parametric estimators of the mean when a scalar outcome is missing at random (e.g., tsiatis2006semiparametric, ; little2019statistical, ; molenberghs2015handbook, ). Popular estimators include outcome regression (also called regression imputation) estimators (e.g., Tan, ; genback2019causal, ) and augmented inverse probability weighted estimators (also called doubly robust estimators; e.g., robins1994estimation, ; bang2005doubly, ; kang2007demystifying, ). The outcome regression (OR) method involves modelling the relationship between the outcome and the observed covariates and imputing the missing outcomes with predictions from this model. The doubly robust (DR) estimator can be interpreted as a correction of the OR estimator, using weights that are based on modelling the missingness mechanism as a function of the covariates. The DR estimator is locally efficient (it attains the semi-parametric efficiency bound if both models are correct) and doubly robust (consistent and asymptotically normal if only one of the two models is correct; (e.g., Seaman:2018, )).
@@
These asymptotic properties allow us to use recent results in Liebl to propose simultaneous confidence bands for the functional mean, providing a given coverage level over the entire domain. At the same time, these bands also allow for local inferences and the interpretation of local significance levels over user-determined partitions of the domain. Up to our knowledge, there are only few earlier peer-reviewed publications on doubly robust estimators in the present context. Belloni , in the causal inference context, study doubly robust estimators and obtain their limiting distribution, but without an expression for the covariance function. Instead, they propose a resampling scheme to obtain valid inference. In another related contribution, ecker2023causal study an OR estimator for a functional causal effect and provides its finite sample distribution, albeit under stricter distributional assumptions about the underlying data.
@@ 2.1 Model and semi-parametric estimation
Suppose we have a random sample, $(\mathcal{Y}_{i},\textbf{x}_{i},Z_{i})_{i=1}^{n}$, from the joint distribution of $(\mathcal{Y},\textbf{x},Z)$, where the outcome variable, $\mathcal{Y}$, is a $L^{2}([0,1],\mathbb{R})$ valued random element, and the set of covariates, x, is a $p\times 1$ random vector with mean vector $\bm{\mu}_{x}$ and covariance matrix $\bm{\Sigma}_{x}$. Further, $Z$ is a binary variable indicating whether the outcome $\mathcal{Y}$ is observed, in which case $Z=1$, or missing ($Z=0$). We assume throughout that $\Pr[Z_{i}=1|\textbf{x}_{i},\mathcal{Y}_{i}]=\Pr[Z_{i}=1|\textbf{x}_{i}]$, that is, the missingness mechanism is ignorable (also referred to as the outcome being missing at random, (tsiatis2006semiparametric, , Chap 6)). The parameter of interest is the functional mean $\mu_{y}=\mathbb{E}\left(\mathcal{Y}\right)$.
@@
$\mathcal{Y}_{i}=\textbf{x}_{i}^{\top}\bm{\beta}+\epsilon_{i},$ (1)
where $\bm{\beta}=(\beta_{1},\dots,\beta_{p})$ and $\beta_{j}$s are non-stochastic functions in $L^{2}([0,1],\mathbb{R})$, and $\epsilon_{i}$ is a $L^{2}([0,1],\mathbb{R})$ valued random element with zero mean and with $\sigma_{\epsilon}$ as its covariance function. Furthermore, a working model for $\Pr[Z_{i}=1|\textbf{x}_{i}]$ is specified as the logistic regression with the conditional probability of being observed in the sample:
$\tau(\textbf{x}_{i}^{\top}\bm{\gamma}),$ (2)
@@
$\hat{\mu}_{OR}=n^{-1}\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\hat{\bm{\beta}}$, (3)
where $\hat{\bm{\beta}}$ is the functional estimation $\left(\mathbf{X}^{\top}\mathbf{X}\right)^{-1}\mathbf{X}^{\top}\bm{\mathcal{Y}},$ with $\bm{\mathcal{Y}}=\left(\mathcal{Y}_{1},\dots,\mathcal{Y}_{n}\right)^{\top}$, and $\mathbf{X}=\left(Z_{1}\mathbf{x}_{1},\dots,Z_{n}\mathbf{x}_{n}\right)^{\top}$. Based on the two working models (1) and (2), the doubly robust estimator of ${\mu}_{y}$, is defined as
@@
where $\hat{\bm{\gamma}}$ is the maximum likelihood estimator of $\bm{\gamma}$.
@@ 2.2 Main results
Our strategy to obtain the asymptotic distribution of the estimators (3) and (4) consists of two steps that allow us to apply Prohorov’s Theorem (see Bill79, , Sect. 5, p.57). First, we consider arbitrary finite-dimensional distributions (i.e. based on a finite discretization on [0,1]) and deduce their joint asymptotic distributions (Lemma 3 and 4 in the Appendix). These results are the consequence of the theory on M-estimators NEWEY19942111 . We provide in these lemmas an explicit characterization of the asymptotic covariance function of multivariate OR and DR estimators, which we cannot find in the earlier literature. In a second step, given the finite-dimensional results, the functional asymptotic results are established by demonstrating tightness. For this purpose, the re-scaled OR estimator, $\sqrt{n}\left(\hat{\mu}_{OR}-\mu_{y}\right)$, can be rewritten as:
$\sqrt{n}\left(\bar{\textbf{x}}_{n}-\bm{\mu}_{x}\right)^{\top}\bm{\beta}+\left(\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\right)\left(\sum_{i=1}^{n}Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right)^{-1}\bm{\zeta}.$ (5)
where $\bm{\zeta}=\left(n^{-1/2}\sum_{i=1}^{n}{Z_{i}x_{i,1}\epsilon_{i}},\dots,n^{-1/2}\sum_{i=1}^{n}{Z_{i}x_{i,p}\epsilon_{i}}\right)^{\top}$. Similarly, the re-scaled DR estimator, $\sqrt{n}\left(\hat{\mu}_{DR}-\mu_{y}\right)$, can be represented as
$\sqrt{n}\left(\hat{\mu}_{OR}-\mu_{y}\right)-\left(\sum_{i=1}^{n}\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\bm{\gamma}})}\textbf{x}_{i}^{\top}\right)\left(\sum_{i=1}^{n}Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right)^{-1}\bm{\zeta}+n^{-1/2}\sum_{i=1}^{n}\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\bm{\gamma}})}\epsilon_{i}.$ (6)
@@
It is worth noting that the functional limiting distribution of the OR estimator (3) could be obtained by applying functional asymptotic theorems developed in recent years, e.g. telschow2022simultaneous . However, the functional DR estimator (4) presents more of a challenge, therefore the strategy outlined above is required. To maintain consistency throughout the paper, this strategy is applied to address both estimators (3) and (4). This allows us to obtain the following results on the asymptotic distributions for these functional estimators.
Theorem 1.
Assume that the working model (1) is correctly specified, i.e., $E(\mathcal{Y}_{i}\mid\mathbf{x}_{i})=\mathbf{x}_{i}^{\top}\bm{\beta}$. Then the functional OR estimator (3) is asymptotically distributed as a Gaussian process:
$\sqrt{n}(\hat{\mu}_{OR}-\mu_{y})\overset{d}{\rightarrow}\mathcal{GP}(0,\bm{\beta}(s)^{\top}\bm{\Sigma}_{\mathbf{x}}\bm{\beta}(t)+\bm{\mu}_{\mathbf{x}}^{\top}\bm{\Pi}^{-1}\bm{\mu}_{\mathbf{x}}\cdot\sigma_{\epsilon}(s,t)).$
@@
By Prohorov’s Theorem, this theorem is a consequence of the results in Lemma 4, 1, and 2; see Appendix A and B for details. ∎
Theorem 2.
Assume that at least one of the working models (1) and (2) is correctly specified, i.e., $E(\mathcal{Y}_{i}\mid\mathbf{x}_{i})=\mathbf{x}_{i}^{\top}\bm{\beta}$ or $\Pr[Z_{i}=1|\mathbf{x}_{i}]=\tau(\mathbf{x}_{i}^{\top}\bm{\gamma})$. Then, the functional DR estimator $\sqrt{n}(\hat{\mu}_{DR}-\mu_{y})$ is asymptotically distributed as a Gaussian process with zero mean function. Further, if both working models are correctly specified, the covariance function simplifies and we have:
$\sqrt{n}(\hat{\mu}_{DR}-\mu_{y})\overset{d}{\rightarrow}\mathcal{GP}(0,\bm{\beta}(s)^{\top}\bm{\Sigma}_{\mathbf{x}}\bm{\beta}(t)+\mathbb{E}[\tau^{-1}(\mathbf{x}_{i}^{\top}\bm{\gamma})]\sigma_{\epsilon}(s,t)).$
@@
By Prohorov’s Theorem, this theorem is a consequence of the results in Lemma 4, 1, and 2; see Appendix A and B for details. ∎
Note that Theorem 2 yields a doubly robustness property for the functional estimator, whereby it is asymptotically distributed as a Gaussian process with zero mean function (concentrates at root-$n$ rate) when only one of the nuisance models is correctly specified. Given that we have functional estimators with Gaussian processes as their limiting distribution, we can obtain simultaneous confidence bands as described below.
@@ 2.3 Feasible estimation and simultaneous confidence bands
The functional estimators (3) and (4) are not directly feasible since the functional outcomes are infinite dimensional objects. However, considering functional estimators and their derived asymptotic distribution allows us to perform simultaneous inference on feasible counterparts. An example of feasible estimators that we use here are the finite-dimensional estimators based on a discretization of the domain in (10) and (11) in Appendix A.
Simultaneous confidence bands for the functional parameter of interest, $\mu_{y}$, can be constructed by applying the results obtained by Liebl . This work introduces ”fast and fair” simultaneous confidence bands based on the construction of an adaptive, non-constant critical value function, which ensures that false positive rates are balanced across different partitions of the domain. The method enables both global and local interpretations of the bands and is less computationally expensive compared to many simulation-based and resampling-based approaches (e.g., Pini17, ; Abram, ).
To apply the results in Liebl , we need to further restrict $\mathcal{Y}_{i}$ to be once continuously differentiable almost surely in the sequel. Under this assumption and the conditions of Theorem 1, a valid 95% simultaneous confidence band (SCB) for $\mu_{y}$ based on the OR estimator is given by:
@@
where $C_{OR}$ is the covariance function of the OR estimator given in Theorem 1, and $\mathit{u}^{\star}_{2.5}(t)$ is a critical value function chosen such that
@@
Simultaneous confidence bands for $\mu_{y}$ based on the DR estimator can be constructed analogously by using $\hat{\mu}_{DR}$ and $C_{DR}$ given in Theorem 2. See Liebl2 ; Liebl for further details on the construction and properties of these simultaneous confidence bands.